# Exact calculation of degrees for lattice equations: a singularity approach

### v2

@@ Abstract.
The theory of degree growth and algebraic entropy plays a crucial role in the field of discrete integrable systems. However, a general method for calculating degree growth for lattice equations (partial difference equations) is not yet known. Here we propose a new method to rigorously compute the exact degree of each iterate for lattice equations. Our strategy is to extend Halburd’s method, which is a novel approach to computing the exact degree of each iterate for mappings (ordinary difference equations) from the singularity structure, to lattice equations. First, we illustrate, without rigorous discussion, how to calculate degrees for lattice equations using the lattice version of Halburd’s method and discuss what problems we need to solve to make the method rigorous. We then provide a framework to ensure that all calculations are accurate and rigorous. We also discuss how to detect the singularity structure of a lattice equation. Our method is not only accurate and rigorous but also can easily be applied to a wide range of lattice equations.
@@ 1. Introduction
The first integrability test for discrete systems in history is singularity confinement [10]. It was proposed as a discrete analogue of the so-called Painlevé property. Roughly speaking, a discrete equation enters a singularity if it loses information on initial values, and an equation is said to pass the singularity confinement test if the lost information is recovered after a finite number of iterations. However, it is widely known today that singularity confinement is not perfect as an integrability test. Not only the so-called Hietarinta-Viallet equation [18] but a large number of non-integrable systems that pass the singularity confinement test have been discovered and constructed. It is also known that most linearizable equations possess non-confined singularities [36].
The most accurate test for discrete integrability today is algebraic entropy. The algebraic entropy of a discrete equation (defined by a rational function) is defined as follows. Each iterate of an equation can be expressed as a rational function in the initial variables and we focus on its degree. Algebraic entropy measures how fast the degree sequence grows as time evolves. Roughly speaking, the algebraic entropy of an equation is 00 (resp. positive) if the degrees grow polynomially (resp. exponentially). If the algebraic entropy of an equation is 00, then the zero algebraic entropy criterion deems the equation integrable. The integrability test by algebraic entropy is so accurate that this criterion is often thought of as defining discrete integrability.
As seen above, the algebraic entropy is defined by degree growth. However, it is very difficult in general to concretely calculate the degrees for an equation. Therefore, developing a method to calculate degree growth is one of the most important problems in the field of discrete integrable systems.
In the case of mappings (ordinary difference equations) of rank $2$, a general theory of degree growth has been established thanks to algebro-geometric methods [9, 5, 7, 42]. Calculating degree growth for rank-$2$ mappings is not difficult if it has good singularity structure, such as to pass the singularity confinement test. For example, if a mapping of rank $2$ passes the singularity confinement test, one can construct an algebraic surface called a space of initial conditions, and the algebraic entropy coincides with the logarithm of the maximum eigenvalue of the linear action on the Picard group. In the case of higher-rank mappings, however, no general classification theory is known yet. How to construct a space of initial conditions and how to calculate degree growth are now being studied [3, 47, 6, 1, 2].
@@
As for degree growth in the case of lattice equations (partial difference equations), a few previous studies are known, such as [43, 46, 19, 12, 31, 40, 41, 20, 13, 14, 21]. For now, however, even rigorously calculating the degree growth for a concrete lattice equation is not easy. In the case of lattice equations, a choice of domain (initial value problem) is essentially important when considering degree growth. According to [20], a domain on which we consider an initial value problem must satisfy some conditions, which we will describe in Definition 3.1), and it is a good strategy to consider the growth of individual degrees (degrees with respect to an arbitrary single initial variable) instead of total degrees (degrees with respect to all initial variables). Therefore, developing a method to calculate individual degrees is an important problem in the field of integrability tests for lattice equations.
@@
It would be natural to think that, now that we have a general classification theory of degree growth in the case of rank-$2$ mappings, the next class of equations we should consider is higher-order mappings, not lattice equations. One of the reasons why lattice equations are much harder to analyze in general than higher-rank mappings is that an initial value problem for a lattice equation has infinitely many initial variables and the degrees of freedom are too large. However, as will be shown in the main part of this paper, our theory is completely based on this fact. Unfortunately, it is not expected that our theory can be applied to the case of mappings.
@@
Let us calculate the degree sequence of the mapping
@@
where $a\neq 0$ is a constant. This equation was first introduced in [37] and the geometric analysis was performed in [32].
@@
  • $(-3a,-a,a,3a)$,
which are both confining and open (non-periodic). That is, for example, if $x_{n-1}$ is generic and $x_{n}=3a+\varepsilon$ where $\varepsilon$ is an infinitesimal parameter, then one obtains
@@
$x_{n+4}$ $=-a+O\left(\varepsilon\right),$
@@
$x_{n+7}$ $=-x_{n-1}+O\left(\varepsilon\right),$
where the exponent of each $O\left(\varepsilon^{\ell}\right)$ above coincides with the leading order of the remaining part. Therefore, the multiplicities of the singularities of this pattern are all $1$.
The equation also has the following periodic patterns of period $2$ and $5$, respectively:
@@
which, exactly speaking, are not singularities since the equation does not lose information on initial values. Note that these patterns extend to both sides. Geometrically speaking, each pattern corresponds to the motion of curves
@@
[…]
First, let us verify all the possibilities where ±$\pm 3a$ or $\infty$ appears as a value. According to the singularity analysis above, we know the following:
  • If $x_{n}(z)$ spontaneously becomes $3a$ for $z=z^{*}\in\mathbb{P}^{1}(\mathbb{C})$, then we have $x_{n+1}(z^{*})=\infty$, $x_{n+3}(z^{*})=\infty$, $x_{n+5}(z^{*})=\infty$, $x_{n+6}(z^{*})=-3a$. That is, if $x_{n-1}(z^{*})$ is generic and $x_{n}(z^{*}+\epsilon)=3a+C\epsilon^{\ell}+O\left(\epsilon^{\ell+1}\right)$ where $\epsilon$ is infinitesimal, $C\neq 0$ and $\ell\in\mathbb{Z}_{>0}$, then we have $x_{n+1}(z^{*}+\epsilon)\sim\epsilon^{-\ell}$, $x_{n+3}(z^{*}+\epsilon)\sim\epsilon^{-\ell}$, $x_{n+5}(z^{*}+\epsilon)\sim\epsilon^{-\ell}$ and $x_{n+6}(z^{*}+\epsilon)+3a\sim\epsilon^{\ell}$. Starting with $z=3a$ is included in this case, too.
  • If $x_{n}$ spontaneously becomes $-3a$ at $z=z^{*}$, then we have $x_{n+3}=3a$.
@@
  • The periodic pattern of period $5$ does not generate ±$\pm 3a$ or $\infty$ as a value.
Let $Z^{+}_{n}$ be the number of $z^{*}\in\mathbb{P}^{1}(\mathbb{C})$ that gives the pattern
@@
counted with multiplicity, i.e.,
$Z^{+}_{n}=\#\left\{z^{*}\in\mathbb{P}^{1}\mid\text{$z=z^{*}$ generate this pattern}\right\}\quad(\text{with multiplicity}).$
For example, the multiplicity of a pattern $(x_{n}=3a,\infty,a,\infty,-a,\infty,-3a)$ is the exponent $\ell$ in the above list. We set $Z^{+}_{n}=0$ ($n\leq 0$) for convenience. Then, using the number of preimages of $\infty$, we obtain
@@
where $\phi^{(2)}_{2m}=1$ and $\phi^{(2)}_{2m+1}=0$ for $m\in\mathbb{Z}$. For example, the term $Z^{+}_{n-1}$ corresponds to the number of $z^{*}$ that gives the pattern $(x_{n-1}=3a,x_{n}=\infty,a,\infty,-a,\infty,-3a)$ and the term $\phi^{(2)}_{n+1}$ corresponds to the number of the periodic pattern $(x_{0}=\text{reg},z=\infty,\text{reg},\infty,\text{reg},\infty,\cdots)$.
In a similar way to the above, using the number of preimages of ±$\pm 3a$, one finds that
$\deg_{z}x_{n}(z)$ $=Z^{+}_{n}+Z^{-}_{n-3}$
$=Z^{+}_{n-6}+Z^{-}_{n},$
where $Z^{-}_{n}$ is the number of patterns $(x_{n}=-3a,-a,a,3a)$ counted with multiplicity (we set $Z^{-}_{n}=0$ for $n\leq 0$).
@@
This system of linear equations with respect to $Z^{\pm}_{n}$ can be thought of as defining the time evolution of $Z^{\pm}_{n}$ with the initial values
@@
$\begin{matrix}n\colon&0,&1,&2,&3,&4,&5,&6,&7,&8,&9,&10,&11,&12,&13,&14,&15,&16,&\cdots,\\ Z^{+}_{n}\colon&0,&1,&1,&2,&2,&3,&4,&5,&6,&7,&8,&10,&11,&13,&14,&16,&18,&\cdots,\\ Z^{-}_{n}\colon&0,&1,&1,&2,&3,&4,&6,&7,&9,&11,&13,&16,&18,&21,&24,&27,&31,&\cdots,\\ \deg_{z}x_{n}(z)\colon&0,&1,&1,&2,&3,&4,&6,&8,&10,&13,&15,&19,&22,&26,&30,&34,&39,&\cdots,\end{matrix}$
@@
One can show in the same way that the degree sequence with respect to the other initial variable, $x_{0}$, grows polynomially as well. Therefore, we conclude that the equation is integrable.
@@
    Let $z$ be an arbitrary one of the initial variables and let the others be generic numerical values.
@@
    Choose some “singular” values $\{\alpha_{1},\ldots,\alpha_{J}\}\subset\mathbb{P}^{1}(\mathbb{C})$ appropriately from the singularity patterns.
@@
    Combining degree relations, obtain a system of linear recurrences for $Z_{n}(\beta_{i})$, which can be thought of as defining the time evolution for $Z_{n}(\beta_{i})$ if the number of degree relations is sufficiently many.
@@
The first problem lies in step 6. As pointed out in [38, Section 5], there are cases where an equation has many singularity patterns but the degree relations do not completely determine a time evolution because the number of linear equations is not sufficient. It is known empirically that such a problem is likely to occur if singularity patterns are short. However, the specific conditions for this phenomenon are not yet known.
@@
In this paper, we construct an easy but rigorous method to calculate exact individual degrees for lattice equations. For simplicity, in this paper, we focus on quad equations
@@
and $\Phi$ depends on all of $B$, $C$ and $D$. Throughout this paper, we always think of the equation as defining the evolution in the northeast direction. We will introduce some conditions on $\Phi$ (see Definitions 3.6 and 3.7) but do not assume other nonessential conditions. For example, we do not assume that the equation can be solved in another direction.
The paper is organized as follows. In §2, we introduce the lattice version of Halburd’s method without rigorous discussion. We will discuss what problems must be solved to make the method rigorous. The main part of this paper is §3 and §4. We will state our main theorems in §3 and give proofs in §4. Since all we need to rigorously calculate degrees are stated in §3, a reader who is not interested in proofs can skip §4. In §5, we analyze several examples to see how to calculate exact degrees with our theory. For a reader who is not familiar with singularity analysis, we discuss how to detect and calculate singularity patterns, too. In Appendix A, we prove Theorem A.1, which states that when considering the degrees for a quad equation, there are only four essential choices of domain. In Appendix B, we recall some definitions and basic properties of substitution and derivative, which play an important role in §4. Appendix C consists of some algebraic lemmas, which we use in §4. In Appendix D, we recall the definition of divisors on $\mathbb{P}^{1}$ and introduce some unusual notations.
@@
The following is the notation we use in this paper. While some are commonly used, many others are introduced for this paper.
  • $\delta_{a,b}$, $\delta^{a,b}_{c,d}=\delta_{a,b}\delta_{c,d}$: Kronecker’s delta.
  • $\overline{K}$: the algebraic closure of a field $K$.
  • $\partial_{X}=\frac{\partial}{\partial X}$: the (partial) derivative with respect to a variable $X$.
  • $\mathbb{P}^{1}(K)=K\cup\{\infty\}$: the projective line over a field $K$.
  • $\operatorname{trdeg}_{K_{1}}K_{2}$: the transcendental degree of a field extension $K_{1}\subset K_{2}$.
@@
  • $O\left(\varepsilon^{\ell}\right)$: the Landau notation, i.e., the leading order of the expansion at $\varepsilon=0$ is $\varepsilon^{\ell}$ or higher.
@@
  • ($\mathbb{Z}^{2}_{\leq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\leq t;\ n^{\prime}\leq n\}$, $H_{\leq(t,n)}=H\cap\mathbb{Z}^{2}_{\leq(t,n)}$: the past light cone emanating from $(t,n)$.
  • ($\mathbb{Z}^{2}_{\geq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\geq t;\ n^{\prime}\geq n\}$, $H_{\geq(t,n)}=H\cap\mathbb{Z}^{2}_{\geq(t,n)}$: the future light cone emanating from $(t,n)$.
  • $\deg_{z}$: the degree of a rational function with respect to $z$.
  • $\mathbb{P}^{1}=\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$: if we simply write $\mathbb{P}^{1}$, the base field is $\overline{\mathbb{K}}$.
  • $\mathbb{K}_{t,n}$: a subfield of $\mathbb{K}$. (Definition 4.2)
[…]
  • $\bigsqcup_{\lambda}T_{\lambda}$: the direct sum of sets.
  • $\deg D$: the degree of a divisor $D$ on $\mathbb{P}^{1}$.
  • $\operatorname{supp}D$: the support of a divisor $D$ on $\mathbb{P}^{1}$.

### v3

@@ Abstract.
The theory of degree growth and algebraic entropy plays a crucial role in the field of discrete integrable systems. However, a general method for calculating degree growth for lattice equations (partial difference equations) is not yet known. Here we propose a method to rigorously compute the exact degree of each iterate for lattice equations. Halburd’s method, which is a novel approach to computing the exact degree of each iterate for mappings (recurrence relations, typically from ordinary difference equations) from the singularity structure, forms the basis of our idea. The strategy is to extend this method to lattice equations. First, we illustrate, without rigorous details, how to calculate degrees for lattice equations using the lattice version of Halburd’s method, and outline the issues that must be resolved to make the method rigorous. We then provide a framework to ensure that all calculations are accurate and rigorous. We further address how to detect the singularity structure in lattice equations. Our method is not only accurate and rigorous but can also be easily applied to a wide range of lattice equations.
@@ 1. Introduction
The first integrability test for discrete systems in history is singularity confinement [11]. It was proposed as a discrete analogue of the so-called Painlevé property. Roughly speaking, a discrete equation enters a singularity if it loses information on initial values, and an equation is said to pass the singularity confinement test if the lost information is recovered after a finite number of iterations. However, it is widely known today that singularity confinement is not perfect as an integrability test. Not only the so-called Hietarinta-Viallet equation [20] but a large number of non-integrable systems that pass the singularity confinement test have been discovered and constructed [4, 41, 12]. It is also known that most linearizable equations possess non-confined singularities [40].
The most accurate test for discrete integrability today is algebraic entropy [5]. The algebraic entropy of a discrete equation (defined by a rational function) is defined as follows. Each iterate of an equation, i.e., each term in the sequence defined by the equation, can be expressed as a rational function in the initial variables and we focus on its degree. Algebraic entropy measures how fast the degree sequence grows as the independent variable tends to infinity. Roughly speaking, the algebraic entropy of an equation is 0 (resp. positive) if the degrees grow polynomially (resp. exponentially). If the algebraic entropy of an equation is 0, then the zero algebraic entropy criterion deems the equation integrable. The integrability test by algebraic entropy shows high empirical accuracy.
As seen above, the algebraic entropy is defined by degree growth. However, it is very difficult in general to compute explicitly the degrees for a given equation. Therefore, developing a method to calculate degree growth is an important problem in the field of discrete integrable systems.
In the case of mappings of rank $2$ (i.e., $2$-dimensional recurrences on a $1$-dimensional lattice, most commonly associated with ordinary difference equations), a general theory of degree growth has been established thanks to algebro-geometric methods [10, 6, 8, 47]. Here, “rank” refers to the number of independent constants in the general solution (degrees of freedom), as commonly used in the theory of ordinary differential and difference equations. Calculating degree growth for a rank-$2$ mapping is not difficult if it has good singularity structure, such as to pass the singularity confinement test. For example, if a mapping of rank $2$ passes the singularity confinement test, one can construct an algebraic surface called a space of initial conditions, and the algebraic entropy coincides with the logarithm of the maximum eigenvalue of the linear action on the Picard group. In the case of higher-rank mappings, however, no general classification theory is known yet. How to construct a space of initial conditions and how to calculate degree growth are now being studied [3, 52, 7, 1, 2].
@@
It would be natural to think that, now that we have a general classification theory of degree growth in the case of rank-$2$ mappings, the next class of equations we should consider is higher-order mappings, not lattice equations. One of the reasons why lattice equations are much harder to analyze in general than higher-rank mappings is that an initial value problem for a lattice equation has infinitely many initial variables. However, as will be shown in the main part of this paper, our theory is based on this infinite nature. Unfortunately, it is not expected that our theory can be applied to the case of mappings (see the discussion in §6).
As for degree growth in the case of lattice equations (partial difference equations), a few previous studies are known, such as [48, 51, 21, 14, 34, 45, 46, 22, 15, 16, 23]. For now, however, even rigorously calculating the degree growth for a concrete lattice equation is not easy. In the case of lattice equations, a choice of domain (initial value problem) is essentially important when considering degree growth. One might think that whether the degree growth of a lattice equation is exponential or polynomial does not depend on the choice of domain. According to [22], however, if considered on a pathological domain, an integrable equation such as the discrete KdV equation has exponential degree growth. The conclusion of [22] is that a domain on which we consider an initial value problem must satisfy some conditions, which we will describe in Definition 3.2, and that it is a good strategy to consider the growth of individual degrees (degrees with respect to an arbitrary single initial variable) instead of total degrees (degrees with respect to all initial variables). For each iterate and every initial variable $z$, we have
[…]
where the sum on the right runs over all initial variables $w$. Under the assumption that the number of the initial variables involved in an iterate increases at most polynomially, the total degree grows at most polynomially if and only if each individual degree does. Therefore, developing a method to calculate individual degrees is an important problem in the study of integrability tests for lattice equations.
@@
Let us calculate the degree sequence of the $3$-point mapping
@@
where $a\neq 0$ is a constant. Note that a $3$-point mapping is a $2$-dimensional ordinary difference equation of the form $x_{n}=F(x_{n-1},x_{n-2})$. This equation was first introduced in [42] and the geometric analysis was performed in [35].
@@
  • (3a,a,a,3a)(-3a,-a,a,3a),
which are both confining and open (not cyclic).
First, we review the definition of singularity confinement on the first pattern. Suppose that $(x_{n-1},x_{n})=(u,3a)$ for generic $u$. Note that the pair $(x_{n-1},x_{n})$ has one degree of freedom since $u$ is generic. The next iterate is $x_{n+1}=\infty$, and the pair $(x_{n},x_{n+1})=(3a,\infty)$ has no degree of freedom. That is, the information on the initial value $u$ is lost, which is the definition of the equation entering a singularity. To calculate further, we introduce an infinitesimal parameter $\varepsilon$ and start with $(x_{n-1},x_{n})=(u,3a+\varepsilon)$ instead. Then, one obtains
@@
(1.1) $x_{n+4}$ $=-a+O\left(\varepsilon\right),$
@@
$x_{n+7}$ $=-u+O\left(\varepsilon\right),$
where the exponent of each $O\left(\varepsilon^{\ell}\right)$ above coincides with the leading order of the remaining part. The important point here is that $x_{n+7}$ is regular, in the sense that it depends on the initial value $u$, and that the pair $(x_{n+6},x_{n+7})=(-3a,-u)$ possesses one degree of freedom. That is, the lost degree of freedom is recovered in a finite number of steps. This is the definition of singularity confinement, and we say that this pattern is confining. On the other hand, a singularity pattern is non-confining if the lost degree of freedom is not recovered. We say an equation is confining (or an equation passes the singularity confinement test) if all singularity patterns of the equation are confining.
Even in the context of singularity confinement, what the term “singularity” exactly denotes is not very clear. In this paper, for a $3$-point mapping, we use this term to denote an iterate of a singularity pattern. More precisely, a singularity $x_{m}=b$ consists of the following data:
  • the value $b$,
  • the position $m$, i.e., the value of the independent variable,
  • the multiplicity, which we will define in the next paragraph.
When we are interested only in the value of a singularity, we use the term “singular value.” For example, the singularity pattern $(3a,\infty,a,\infty,-a,\infty,-3a)$ consists of $7$ singularities but has $5$ singular values: $\pm 3a,\pm a,\infty$. We will define a singularity for lattice equations in the main part of the paper (Definition 3.5).
Let us review the definition of the multiplicity of a singularity. Consider a singularity $x_{m}=b$ for $b\in\mathbb{P}^{1}(\mathbb{C})=\mathbb{C}\cup\{\infty\}$. If $b\in\mathbb{C}$, the expansion of $x_{m}$ around $\varepsilon=0$ is
$x_{m}=b+f(u)\varepsilon^{r}+O\left(\varepsilon^{r+1}\right)$
for some $f(u)\neq 0$ and $r>0$. On the other hand, if $b=\infty$, then the expansion is
$x_{m}=f(u)\varepsilon^{-r}+O\left(\varepsilon^{-r+1}\right)$
for some $f(u)\neq 0$ and $r>0$. In both cases, we define the multiplicity of the singularity $x_{m}=b$ as $r$. Note that this definition coincides with the definition of the multiplicity of $x_{m}=b$ (as a rational function in $\varepsilon$) at $\varepsilon=0$ (see Definition D.1). For example, it follows from the calculation from $x_{n+1}$ to $x_{n+7}$ that the multiplicities of the singularities of the pattern $(3a,\infty,a,\infty,-a,\infty,-3a)$ are all $1$.
Note that the first pattern does not continue with the second one since the next iterate after $(3a,\infty,a,\infty,-a,\infty,-3a)$ is regular, i.e., it depends on $u$ (see the calculation of $x_{n+7}$ in (1.1)). It is known (but does not directly follow from the calculation (1.1)) that the pattern $(3a,\infty,a,\infty,-a,\infty,-3a)$ has no other singularities [35]. In particular, this pattern is not cyclic, in the sense that the pair $(g(u),3a)$ with non-constant function $g(u)$ does not appear afterward. A confining singularity pattern is defined to be open if it has this non-cyclic property.
It is known that the singularity pattern (3a,a,a,3a)(-3a,-a,a,3a) is also confining and open [35]. Calculating the Jacobian of the map $(x_{n-1},x_{n})\mapsto(x_{n},x_{n+1})$, one obtains that the equation has no other singularity pattern. Therefore, this equation passes the singularity confinement test.
To compute the degree sequence by Halburd’s method, we need to know when the equation takes singular values. Searching for the possibility of the equation taking the values $\pm a$, $\pm 3a$, $\infty$, one obtains that the equation has the following periodic patterns of period $2$ and $5$, respectively:
@@
where “reg” denotes some regular value. It should be stressed that these patterns are not singularity patterns since the equation does not lose the dependence on the initial value. For example, we obtain the first pattern by starting from the initial value $(x_{n-1},x_{n})=\left(u,\infty\right)$ with generic $u$. Iterating the recurrence, we have
$x_{n+1}=\infty,\quad x_{n+2}=f_{2}(u),\quad x_{n+3}=\infty,\quad x_{n+4}=f_{4}(u),\quad\cdots,$
where each $f_{j}(u)$ is a non-constant function. Observe that in this example, $f_{j}(u)$ coincides with $u$, though this is not essential. What is important here is that each pair $(x_{m},x_{m+1})$ has one degree of freedom and that the dependence on the initial value $u$ is preserved throughout. Therefore, this pattern is not a singularity pattern. In the second pattern, each pair contains at least one regular value. Therefore, this pattern is not a singularity pattern, either.
A direct calculation shows that these two periodic patterns extend to both sides. Geometrically speaking, each pattern corresponds to the motion of curves
@@
The key idea of Halburd’s method is to calculate $\deg_{z}x_{n}(z)$ as the number of preimages of $x_{n}=\text{(constant)}$ counted with multiplicity for various singular values. It is well known that the degree of a rational function $f\colon\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$, which is defined as the maximum of the degrees of its numerator and denominator when written in lowest terms, coincides with the number of the preimages $f^{-1}(\alpha)$ counted with multiplicity for any $\alpha\in\mathbb{P}^{1}(\mathbb{C})$ (Proposition D.6). This property is based on the fact that any degree-$m$ polynomial over an algebraically closed field has exactly $m$ roots, counted with multiplicities.
First, let us verify all the possibilities where $\pm 3a$ or $\infty$ appears as a value. Since the term “spontaneously” plays a key role in what follows but lacks a precise definition in the literature, we clarify its meaning here: for $z^{*},\alpha\in\mathbb{P}^{1}(\mathbb{C})$, we say that $x_{n}(z^{*})$ spontaneously becomes $\alpha$ if $x_{0}(z^{*}),\ldots,x_{n-1}(z^{*})$ are generic and $x_{n}(z^{*})=\alpha$. This interpretation is adopted only for the present discussion.
According to the singularity analysis above, we know the following:
  • If $x_{n}(z)$ spontaneously becomes $3a$ for $z=z^{*}\in\mathbb{P}^{1}(\mathbb{C})$, then we have $x_{n+1}(z^{*})=\infty$, $x_{n+3}(z^{*})=\infty$, $x_{n+5}(z^{*})=\infty$, $x_{n+6}(z^{*})=-3a$. That is, if $x_{n-1}(z^{*})$ is generic and $x_{n}(z^{*}+\varepsilon)=3a+C\varepsilon^{\ell}+O\left(\varepsilon^{\ell+1}\right)$ where $\varepsilon$ is infinitesimal, $C\neq 0$ and $\ell\in\mathbb{Z}_{>0}$, then we have $x_{n+1}(z^{*}+\varepsilon)\sim\varepsilon^{-\ell}$, $x_{n+3}(z^{*}+\varepsilon)\sim\varepsilon^{-\ell}$, $x_{n+5}(z^{*}+\varepsilon)\sim\varepsilon^{-\ell}$ and $x_{n+6}(z^{*}+\varepsilon)+3a\sim\varepsilon^{\ell}$. Starting with $z=3a$ is included in this case, too.
  • If $x_{n}$ spontaneously becomes 3a-3a at $z=z^{*}$, then we have $x_{n+3}=3a$.
@@
  • The periodic pattern of period $5$ does not generate $\pm 3a$ or $\infty$ as a value.
We use the notation
@@
to denote the singularity pattern $(3a,\infty,a,\infty,-a,\infty,-3a)$ starting at index $n$. Let $Z^{+}_{n}$ be the number of $z^{*}\in\mathbb{P}^{1}(\mathbb{C})$ that gives the pattern $(x_{n}=3a,\infty,a,\infty,-a,\infty,-3a)$ counted with multiplicity, i.e.,
$Z^{+}_{n}=\#\left\{z^{*}\in\mathbb{P}^{1}(\mathbb{C})\mid\text{$z=z^{*}$ generate this pattern}\right\}\quad(\text{with multiplicity}).$
Here, if the first singularity of a pattern is $x_{n}(z^{*})=\alpha\in\mathbb{P}^{1}(\mathbb{C})$, then we define the multiplicity of the pattern as the multiplicity of $x_{n}=\alpha$ at $z=z^{*}$. For example, the multiplicity of the pattern $(x_{n}=3a,\infty,a,\infty,-a,\infty,-3a)$ is the exponent $\ell$ in the above list. We set $Z^{+}_{n}=0$ ($n\leq 0$) for convenience. Then, using the number of preimages of $\infty$, we obtain
@@
where $\phi^{(2)}_{2m}=1$ and $\phi^{(2)}_{2m+1}=0$ for $m\in\mathbb{Z}$. For example, the term $Z^{+}_{n-1}$ corresponds to the number of $z^{*}$ that gives the pattern $(x_{n-1}=3a,\infty,a,\infty,-a,\infty,-3a)$ and the term $\phi^{(2)}_{n+1}$ corresponds to the number of the periodic pattern $(x_{0}=\text{reg},z=\infty,\text{reg},\infty,\text{reg},\infty,\cdots)$.
In a similar way to the above, using the number of preimages of $\pm 3a$, one finds that
$\deg_{z}x_{n}(z)$ $=Z^{+}_{n}+Z^{-}_{n-3}$ $(\text{$\#$ of preimages of $3a$})$
$=Z^{+}_{n-6}+Z^{-}_{n}$ $(\text{$\#$ of preimages of $-3a$}),$
where $Z^{-}_{n}$ is the number of $z^{*}$ that gives the pattern $(x_{n}=-3a,-a,a,3a)$ counted with multiplicity (we set $Z^{-}_{n}=0$ for $n\leq 0$).
@@
This system of linear equations with respect to $Z^{\pm}_{n}$ can be thought of as defining the evolution of $Z^{\pm}_{n}$ with the initial values
@@
[…]
@@
One can show in the same way that the degree sequence with respect to the other initial variable, $x_{0}$, grows polynomially as well. Therefore, the algebraic entropy of this equation is 0, and the integrability criterion by degree growth deems the equation integrable.
@@
    Let $z$ be one of the initial variables and let the others be generic numerical values.
@@
    Choose some singular values $\{\alpha_{1},\ldots,\alpha_{J}\}\subset\mathbb{P}^{1}(\mathbb{C})$ appropriately from the singularity patterns.
@@
    Combining degree relations, obtain a system of linear recurrences for $Z_{n}(\beta_{i})$, which can be thought of as defining the evolution for $Z_{n}(\beta_{i})$ if the number of degree relations is sufficiently many.
@@
The first problem lies in step 6. As pointed out in [43, Section 5], there are cases where an equation has many singularity patterns but the degree relations do not determine an evolution because the number of linear equations is not sufficient. It is known empirically that such a problem is likely to occur if singularity patterns are short. However, the specific conditions for this phenomenon are not yet known.
@@
In this paper, we construct an easy but rigorous method to compute exact individual degrees for lattice equations. For simplicity, in this paper, we focus on quad equations
@@
and $\Phi$ depends on all of $B$, $C$ and $D$. Throughout this paper, we always think of the equation as defining the evolution in the northeast direction. We will introduce some conditions on $\Phi$ (see Definitions 3.9 and 3.11) but do not assume other nonessential conditions. For example, we do not assume that the equation can be solved in another direction.
The paper is organized as follows. In §2, we introduce the lattice version of Halburd’s method without rigorous discussion. We will discuss what problems must be solved to make the method rigorous. The main part of this paper is §3 and §4. We will state our main theorems in §3 and give proofs in §4. Since all we need to rigorously calculate degrees are stated in §3, a reader who is not interested in proofs can skip §4. In §5, we analyze several examples to see how to calculate exact degrees with our theory. For a reader who is not familiar with singularity analysis, we discuss how to detect and calculate singularity patterns, too. In Appendix A, we prove Theorem A.1, which states that when considering the degrees for a quad equation, there are only four essential choices of domain. In Appendix B, we recall some definitions and basic properties of substitution and derivative, which play an important role in §4. Appendix C consists of some algebraic lemmas, which we use in §4. In Appendix D, we recall the definition of divisors on $\mathbb{P}^{1}(K)$ and introduce some unusual notations.
@@
We summarize the notation that will be used in the main part of the paper. For clarity, we separate it into three groups. Notation that appears only in the introduction, is not important for understanding the main results, or occurs only within proofs is explained in the text and not included here.
  1. (1)
    Fundamental concepts and notation (excluding very basic symbols):
@@
    • $\operatorname{div}(f=\alpha\mid T)$, $\operatorname{div}\left(f=\alpha\mid\text{(condition)}\right)$: a divisor on $\mathbb{P}^{1}(K)$. (Definition D.3)
    • $T_{1}\sqcup T_{2}$, $\bigsqcup_{\lambda}T_{\lambda}$: the direct sum of sets (disjoint union).
  2. (2)
    Notation for the setting (main symbols only):
@@
  3. (3)
    Concepts and terminology used in this paper (key concepts only):
    • reg: a regular value, i.e. a value that depends on some initial value.
    • Degree Relation: the expression of the degree of an iterate as the number of preimages of a $\mathbb{P}^{1}(\mathbb{C})$-value.
    • $\mathbb{Z}^{2}_{\leq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\leq t;\ n^{\prime}\leq n\}$, $H_{\leq(t,n)}=H\cap\mathbb{Z}^{2}_{\leq(t,n)}$: the past light cone emanating from $(t,n)$.
    • $\mathbb{Z}^{2}_{\geq(t,n)}=\{(t^{\prime},n^{\prime})\in\mathbb{Z}^{2}\mid t^{\prime}\geq t;\ n^{\prime}\geq n\}$, $H_{\geq(t,n)}=H\cap\mathbb{Z}^{2}_{\geq(t,n)}$: the future light cone emanating from $(t,n)$.
    • Constant Singularity, Constant Singularity Pattern, Singular Value: Definition 3.5.
    • Movable, Fixed, Confining, and Solitary Patterns: Definition 3.5.
    • Basic Pattern: Definition 3.7.
@@
    • $\partial$-factor condition: Definition 3.9.
    • Basic Pattern Condition: Definition 3.11.
    • First Singularity, Starting Value: Definition 3.15.
    • Spontaneous Occurrence of $x_{t,n}(z^{*})=\alpha$: Definition 4.14.
    • $Z_{s,m}(\beta_{i})$: the number of spontaneous occurrences of $x_{s,m}(z)=\beta_{i}$ with multiplicity. (Definition 4.19)